Mary Smith was instructed to drink the following prep for her x-ray:
2 1/2 oz every 15 minutes for 2 hours
4 oz every 30 mins for 2 hours
6 1/2 oz every hour for 2 hours
What is the total quanity of prep you need to give Mary Smith?

Answer by ewatrrr(24785)


Hi
2 1/2 oz every 15 minutes for 2 hours = 5/2oz*8 = 20oz
4 oz every 30 mins for 2 hours = 4oz*4 = 16oz
6 1/2 oz every hour for 2 hours = 13/2*2 = 13oz